OpenStudy (anonymous):

(x^2/9)/(x^1/6) and (g^3/4)/(g^2/3)

OpenStudy (anonymous):

not add, subract!

OpenStudy (anonymous):

so it's simply x^1/3

OpenStudy (anonymous):

\[\frac{2}{9}-\frac{1}{6}=\frac{1}{18}\]

OpenStudy (anonymous):

no answer is \[x^{\frac{1}{18}}\]

OpenStudy (anonymous):

you have to subtract the fraction by finding common denominator etc, not by subtracting top and bottom yes?

OpenStudy (anonymous):

oh ok got it